Given m=20kg,s=3m,f=150N,angle=30degree and myu=0.3 find work


1
Expert's answer
2019-09-27T09:57:49-0400

If we need to find the work done by someone who applies the force of 150N, it will be


"A_F=Fs\text{ cos}\alpha=150\cdot3\cdot\text{cos}30^\circ=389.71\text{ J}."

However, if we are to determine the work done by friction, it will be


"A_f=fs=\mu Ns=\mu(mg-F\space\text{sin}\alpha)s=108.9\text{ J},"

since the force has a component oriented upward while the force of gravity is directed downward, and the force of friction is defined by the normal force, i.e. the difference between the force of gravity and the vertical component of the force of 150 N.